OpenStudy (anonymous):

Please Solve and EXPLAIN!! –y + 3x = 6 y = –6x + 12

Parth (parthkohli):

You can solve by elimination. Where is the problem?

OpenStudy (anonymous):

that's a simultaneous equation, yes?

OpenStudy (anonymous):

yes

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow y + 6x = 12 }\) Now, multiply the first equation by 2 and then subtract both.

Parth (parthkohli):

Or, substitution would also work.

Parth (parthkohli):

If y = -6x + 12, then you can substitute -y with -(-6x + 12)

OpenStudy (anonymous):

-y+3x=6_y=-6x+12 Since 3x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3x from both sides. -y=-3x+6_y=-6x+12 Divide each term in the equation by -1. y=3x-6 y=-6x+12 Since the equation was solved for y, replace all occurrences of y in the other equations with the solution (3x-6). y=3x-6 Move all terms not containing x to the right-hand side of the equation. y=3x-6 9x-6=12 Move all terms not containing x to the right-hand side of the equation. y=3x-6 9x=18 Since the equation was solved for x, replace all occurrences of x in the other equations with the solution (2). y=3x-6 x=2 Since the equation was solved for x, replace all occurrences of x in the other equations with the solution (2). y=3(2)-6x=2 Multiply 3 by each term inside the parentheses. y=6-6 x=2 Subtract 6 from 6 to get 0. y=0 x=2 This is the solution to the system of equations. y=0 x=2

OpenStudy (anonymous):

6x-12+3x=6 9x=18 x=2 –y + 3(2) = 6 y=0 by substitution

OpenStudy (anonymous):

kk